OpenStudy (anonymous):

show all your work solve 2x=14

OpenStudy (anonymous):

14 / 2 = 7

OpenStudy (anonymous):

x = 7

OpenStudy (anonymous):

yah lol

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 2x=14 }\) or \(\large\color{black}{ \displaystyle 2^x=14 }\) ?

OpenStudy (anonymous):

the frist one

OpenStudy (anonymous):

first

OpenStudy (solomonzelman):

even better, although for the second one: \(\large\color{black}{ \displaystyle 2^x=14 }\) \(\large\color{black}{ \displaystyle \color{red}{\log_2} 2^x=\color{red}{\log_2}14 }\) \(\large\color{black}{ \displaystyle x{\tiny~}\log_2 2=\color{re}{\log_2}14 }\) \(\large\color{black}{ \displaystyle x{\tiny~}=\color{rd}{\log_2}14 }\)

OpenStudy (anonymous):

so u juat dived 14 by 2

OpenStudy (anonymous):

yes